# Unlock the Power of Frozen LLMs in Knowledge Graph Completion

### v1

@@ Abstract
Classical knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs). Large Language Models (LLMs) learn extensive knowledge from large corpora with powerful context modeling, which is ideal for mitigating the limitations of previous methods. Directly fine-tuning LLMs offers great capability but comes at the cost of huge time and memory consumption, while utilizing frozen LLMs yields suboptimal results. In this work, we aim to leverage LLMs for KGC effectively and efficiently. We capture the context-aware hidden states of knowledge triples by employing prompts to stimulate the intermediate layers of LLMs. We then train a data-efficient classifier on these hidden states to harness the inherent capabilities of frozen LLMs in KGC. We also generate entity descriptions with subgraph sampling on KGs, reducing the ambiguity of triplets and enriching the knowledge representation. Extensive experiments on standard benchmarks showcase the efficiency and effectiveness of our approach. We outperform classical KGC methods on most datasets and match the performance of fine-tuned LLMs. Additionally, compared to fine-tuned LLMs, we boost GPU memory efficiency by 188×188$\times$188 × and speed up training+inference by 13.48×13.48$\times$13.48 ×.
@@ 1 Introduction
Knowledge graph completion has become a crucial endeavor for the effective application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b). Classical knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which suffer from the sparseness of KGs (Yao et al., 2019). Leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs can mitigate this challenge by utilizing the rich semantics within knowledge graphs, thereby enhancing the performance of knowledge graph completion (Wei et al., 2023).
@@
These approaches generally fall into two types: using frozen LLMs and employing fine-tuned LLMs. The first type leverages frozen language models (e.g., LLaMA and GPT-4) with prompt techniques, such as in-context learning, which facilitate the completion of knowledge graphs directly in natural language outputs (Wei et al., 2023; Yao et al., 2023), offering an efficient solution that utilizes the pre-trained capabilities of these models without the need for further training. The second type involves fine-tuning the language models using a range of supervised techniques, including instruction tuning, to tailor the models specifically for KGC tasks. This method involves fine-tuning language models, such as LLaMA and T5, through the use of training samples (Saxena et al., 2022; Yao et al., 2023). Although this can lead to better performance (Zhang et al., 2023; Yao et al., 2023), it also requires significant increases in time and resource expenditure. A natural question is: How can we utilize LLMs to accomplish KGC tasks with both efficacy and efficiency?
@@
To answer this question, we build upon the efficiency of frozen LLMs and explore the reasons behind the underperformance of frozen LLMs compared to their fine-tuned counterparts. First, while LLMs contain vast amounts of world knowledge encoded within their parameters, accessing the knowledge and decoding it into responses effectively through frozen models can be challenging (Wang et al., 2023). It is well known that LLMs suffer from hallucinations (Zhao et al., 2023), which makes them unreliable for achieving high performance in KGC tasks. Fine-tuning directly modifies the models’ weights through the training samples to improve the retrieval and application of relevant world knowledge, thus leading to better performance (Zhou et al., 2024). Second, KGs often involve complex relationships and ambiguous entity representations (Li et al., 2024), which can be difficult for a frozen LLM to handle accurately without specific tuning. Fine-tuned models can be explicitly trained to interpret and disambiguate such complexities on the basis of the training data provided, linking entities in the KG with the LLM.
@@
To overcome the limitations of frozen LLMs in KGC tasks, we have introduced probing techniques utilizing stimulation prompts, complemented by detailed descriptions of entities, as shown in Figure 1. These methods are designed to effectively extract and leverage the latent knowledge embedded within the models’ parameters, and minimize ambiguous entity representations, thereby enhancing its application in KGC without the need for extensive retraining. Specifically, our research concentrates on the core issues of knowledge graph completion, which are triple classification and relation prediction. We stimulate the intermediate layers of LLMs to capture the context-aware hidden states of knowledge triples. Moreover, we improve the precision of knowledge representation by aligning KG entities with the LLMs, reducing the ambiguity of knowledge triples through the generation of entity descriptions via subgraph sampling from KGs. We train a data-efficient classifier on these obtained hidden states suppervised by the labels, to fully harness the inherent KGC capabilities of the frozen LLMs.
@@
  • We introduce a novel probing method tailored for KGC tasks, designed to stimulate language models to verify knowledge graph triples directly. We further generate knowledge graph entity descriptions through the use of subgraphs, which are then seamlessly integrated into our probing techniques.
  • Our approach employs LLMs with both effectiveness and efficiency to achieve KGC. By utilizing only a subset of the dataset, our method achieves a relative performance improvement of up to 2.7% compared to fine-tuning the LLM on the entire training set. Additionally, it significantly reduces GPU memory usage by a factor of 188 during the training phase and descreases time consumption by a factor of 13.48, all while maintaining comparable performance. [^1] For a detailed comparison of efficiency, please refer to Table 8 in the Appendix B.3.
@@ 2 Method
Our method’s design stems from an analysis of the two primary reasons for the underperformance of frozen LLMs compared to fine-tuned counterparts. First, LLM may not always trigger the depth of knowledge needed and reveal the right responses. We demonstrate in Section 2.2 how to achieve this for the frozen LLM using probing techniques. Second, KGs often involve ambiguous entity representations, which can be difficult for LLM to handle accurately. In Section 2.3 we show how to achieve this with the generation of entity descriptions.
@@ 2.1 Formulation and Notations
A knowledge graph can be represented as a set of triples $G=\{(h,r,t)\}$, where $E$ and $R$ denote the set of entities and relations in G, respectively, $h\in E$ is a head entity, $t\in E$ is a tail entity, we collectively refer to the head entity and tail entity as $e$, and $r\in R$ represents the relation between them. $D$ is the set of descriptions for each entity and relation. We denote $D(e)$, $D(r)$ as the textual description of each entity and each relation.
@@
2.2 Stimulating the Intermediate Layers
Knowledge graphs consist of numerous triples, and the format of these triples is challenging for the LLM to understand, which may mislead the LLM to retrieve irrelevant knowledge and reveal the wrong responses. To minimize the influence of task-unrelated information and enhance the performance of LLMs in KGC, we leverage the internal judgment mechanisms of the language model. Our method involves strategically triggering the necessary depth of knowledge embedded within the LLM to elicit the correct responses required for accurate KGC. By focusing on extracting classification results directly from the model’s internal processing, we ensure that the responses are both relevant and precise, thereby improving the effectiveness of the KGC task.
Since different layers of the language model have different roles (Azaria and Mitchell, 2023), we determine the most suitable layer for the task by experimenting with the validation set. First, we construct prompt templates to stimulate the model to generate the required classification information in the intermediate layers. In the triple classification, our goal is to determine whether the triples in the knowledge graph are true. Therefore, we use the following prompt template, $PT_{1}$, to create positive and negative sample pairs, as shown in Figure 2. Notice that the prompt does not describe positive or negative information for triples.
@@ 2.2 Stimulating the Intermediate Layers
$v\left(s^{+}\right)=\{M_{1-i}\left(s^{+}\right)[-1],0<i<L\},$ (1)
where $M$ represents the language model, $M_{1-i}$ represents the first $i$ layers of the model, $[-1]$ refers to the hidden state corresponding to the last token, $L$ denotes the number of layers in the model. The dimension of $M_{1-i}(s^{+})[-1]$ is the embedding dimension of the language model. Negative samples follow the same principle.
At last, we train a classification model using $v(s^{+})$ and $v(s^{-})$ to capture plausibility information. The goal is to learn a boundary that can separate the hidden states of the intermediate layers of the language model corresponding to positive samples (which are true) from those corresponding to negative samples (which are false). We can use classification models such as multilayer perceptron (MLP), SVM, etc., for classification training. Taking logistic regression to determine the plausibility of samples as an example:
@@
where x represents $v(s)[l]$, $l$ is the layer index, $w$ and $b$ are weights and bias to be learned. For relation prediction, we simply transform the binary classification into multi-class classification. By utilizing the prompt in Figure 7, we stimulate the LLM to generate hidden states that capture the characteristics of different relation types in the KG. Details can be found in the Appendix B.1.
The task-unrelated information does not affect the classification results. The reason is that the positive samples $S^{+}$ and negative samples $S^{-}$ are constructed using the same prompt template. The classifier only needs to capture the plausibility of the triples in $v(S^{+})$ and $v(S^{-})$, that is, whether the triple is true or not. The advantage of using this method for KGC is that it can ignore the impact of the task-unrelated information and reveal the right responses for accurate KGC.
@@ 2.3 Subgraph Entity Description Generator
The semantics of entities in the knowledge graph are highly complex; it is difficult for the language model to understand the entity based solely on the entity’s name. Therefore, we inform the language model about the meaning of each entity in the input. The meaning of each entity in KG is generated by its one-hop subgraph.
@@
To address incomplete entity descriptions within certain datasets, we have devised a method to generate $D(e)$. We construct $D(e)$ in two ways. The first approach involves a straightforward method where we concatenate all the triples within the one-hop subgraph surrounding the entity $e$, thereby forming $D(e)$. However, this direct concatenation can sometimes lead to ineffective or misleading stimuli for the LLM, as the structure and content of these subgraph triples may significantly differ from the text seen during the LLM’s pre-training phase. This divergence can adversely affect the model’s response accuracy. To mitigate this issue, our second approach leverages a language model to transform these triples into a format that is more comprehensible to the LLM. By rephrasing the triples from the one-hop subgraph around the entity $e$ into natural language text, we produce a more refined and model-friendly version of $D(e)$:
@@
where $\operatorname{subgraph(\cdot)}$ is a function that retrieves the set of all triples in the one-hop subgraph where the entity resides; $\operatorname{transform(\cdot)}$ is a function that converts the triple into sentence; $\operatorname{concat(\cdot)}$ is a function to process triple statements. For the first approach, $\operatorname{concat(\cdot)}$ is simply a string concatenation function that joins each transformed triple statement with a separator. For the second approach, we need to utilize LLM with the in-context learning technique (Brown et al., 2020) to generate descriptions of entities based on the triples in their subgraph. Please refer to Appendix A.1 for details.
@@ 3 Experiments
  • RQ3: How about data efficiency and compute efficiency of our proposed method?
@@ 3.2 Experimental Settings
LLAMA-7B N-L 0 0.091 0.211 0.573 0.458 0.658
ALPACA-7B N-L 0 0.226 0.696 0.561 0.714 0.526
LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
ALPACA-7B-ICL N-L 2 0.675 0.706 0.607 0.691 0.605
ALPACA-DES N-L 0 0.611 - 0.590 0.548 0.525
LLAMA-MLP H-MLP 10,000 0.851 0.874 0.679 0.866 0.882
LLAMA-DES-H H-MLP 10,000 0.895 - 0.736 0.935 0.862
LLAMA-DES2 (Tri) H-MLP 10,000 0.847 0.764 0.673 0.785 0.782
LLAMA-DES2 (GPT)* H-MLP 1,000 0.890 0.892 0.705 0.907 0.783
Table 2: The main experiment results (accuracy) of triple classification. In the ’Pred’ column, S-F uses score function, L-MLP uses MLP on last layer hidden states, N-L uses natural language, and H-MLP uses MLP on intermediate layer hidden states to predict. ’Samples’ indicates training sample counts. ALPACA-DES combines non-generated descriptions with $PT_{2}$ for Alpaca-7B; LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimuli generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimuli generation. Top three results per dataset are bolded.
@@ 3.3 Main Results (For RQ1)
From Table 2, we can observe that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples. Additionally, LLAMA-DES-H and LLAMA-DES2 (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3. Moreover, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B, ALPACA-7B, LLAMA-7B-ICL and ALPACA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2. However, it is noteworthy that ALPACA-DES fails to outperform ALPACA-7B, suggesting that mere incorporation of entity descriptions may not yield direct performance enhancements, considering the possibility of LLMs revealing the wrong responses and being influenced by hallucinations. Thus, the experimental results substantiate the effectiveness of our proposed methodology. Additionally, the frozen language model can also achieve similar KGC effects as fine-tuning the language model.
Notice that the performance on the UMLS dataset is special, where using generated descriptions impedes the classification accuracy of the model, and using the non-generated entity descriptions is also worse than not using any entity descriptions. It is postulated that the underlying reason may lie in the specialized nature of entity names in the UMLS dataset, resulting in minimal ambiguity. The entities in the knowledge graph have well-aligned with the internal entities of the LLM. Meanwhile, using entity descriptions may be affected by uncleaned description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states.

### v2

@@ Abstract
Traditional knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs). Large Language Models (LLMs) learn extensive knowledge from large corpora with powerful context modeling, making them promising for mitigating the limitations of previous methods. Directly fine-tuning LLMs offers great capability but comes at the cost of huge time and memory consumption, while utilizing frozen LLMs yields suboptimal results. In this work, we aim to leverage LLMs for KGC effectively and efficiently. We capture the context-aware hidden states of knowledge triples by employing prompts to stimulate the intermediate layers of LLMs. We then train a data-efficient classifier on these hidden states to harness the inherent capabilities of frozen LLMs in KGC. Additionally, to reduce ambiguity and enrich knowledge representation, we generate detailed entity descriptions through subgraph sampling on KGs. Extensive experiments on standard benchmarks demonstrate the efficiency and effectiveness of our approach. We outperform traditional KGC methods across most datasets and, notably, achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by 188×188$\times$188 × and accelerating training and inference by 13.48×13.48$\times$13.48 ×.
Unlock the Power of Frozen LLMs in Knowledge Graph Completion
Bo Xue, Yi Xu, Yunchong Song, Yiming Pang, Yuyang Ren, Jiaxin Dingthanks:  Jiaxin Ding is the corresponding authors., Luoyi Fu, Xinbing Wang Shanghai Jiao Tong University, Shanghai, China {sappho_x, jiaxinding, yiluofu} @sjtu.edu.cn
@@ 1 Introduction
Knowledge graph completion has become a crucial endeavor to improve the application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b). Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which suffer from the sparseness of KGs (Yao et al., 2019). Leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs can mitigate this challenge by utilizing the rich semantics within knowledge graphs, thereby enhancing the performance of knowledge graph completion (Wei et al., 2023).
@@
These approaches generally fall into two types: using frozen LLMs with prompting techniques and fine-tuned LLMs specifically for KGC. The first type leverages frozen LLMs with prompt techniques, such as in-context learning. These techniques enable models to perform KGC directly through natural language outputs (Wei et al., 2023; Yao et al., 2023), harnessing the pre-trained knowledge embedded within these models without additional training. While this approach offers computational efficiency, it may not fully capture the task-specific nuances required for optimal KGC performance. The second type involves fine-tuning the LLMs using supervised techniques, including instruction tuning, to adapt models specifically for KGC tasks. This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2023), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2023; Yao et al., 2023). However, fine-tuning demands significant increases in time and resource expenditure. This dichotomy raises a fundamental question: How can we utilize LLMs to accomplish KGC tasks with both efficacy and efficiency?
@@
To answer this question, we explore the reasons behind the underperformance of frozen LLMs compared to their fine-tuned counterparts. First, while LLMs contain vast amounts of world knowledge encoded within their parameters, accessing the knowledge and decoding it into responses effectively through frozen models can be challenging (Wang et al., 2023). LLMs can produce hallucinations-plausible but incorrect outputs (Zhao et al., 2023)-making them unreliable for precise KGC tasks. Fine-tuning directly modifies the models’ weights through task-specific data enhancing the retrieval of relevant knowledge and improving performance (Zhou et al., 2024). Second, KGs often involve complex relationships and ambiguous entity representations (Li et al., 2024), which can be difficult for a frozen LLM to handle such nuances accurately without specific tuning. Fine-tuned models can be explicitly trained to interpret and disambiguate the complexities enabling better linkage of entities within KGs.
@@
To overcome the limitations of frozen LLMs in KGC tasks, we introduce probing techniques that utilize stimulation prompts, complemented by detailed entity descriptions, as shown in Figure 1. These methods are designed to effectively extract and leverage the latent knowledge embedded within the models’ parameters, and minimize ambiguous entity representations, thereby enhancing its application in KGC without the need for extensive retraining. Specifically, our research concentrates on the core tasks of knowledge graph completion: triple classification and relation prediction.
We stimulate the intermediate layers of LLMs to capture the context-aware hidden states of knowledge triples. Moreover, we improve the precision of knowledge representation by aligning KG entities with the LLMs, reducing the ambiguity of knowledge triples through the generation of entity descriptions via subgraph sampling from KGs. We train a data-efficient classifier on the obtained hidden states suppervised by labels, to fully harness the inherent KGC capabilities of the frozen LLMs.
@@
  • We introduce a novel probing method tailored for KGC tasks, designed to stimulate language models to predict knowledge graph triples directly. We further generate knowledge graph entity descriptions through the use of subgraphs, which are then seamlessly integrated into our probing techniques.
  • Our approach employs LLMs with both effectiveness and efficiency to achieve KGC. By utilizing only a subset of the dataset, our method achieves a relative performance improvement of up to 2.7% compared to fine-tuning the LLM on the entire training set. Additionally, it significantly reduces GPU memory usage by a factor of 188 during the training phase and decreases time consumption by a factor of 13.48, all while maintaining comparable performance. [^1] For a detailed comparison of efficiency, please refer to Table 8 in the Appendix B.3.
@@ 2 Method
Our method is designed based on an analysis of the two primary reasons for the underperformance of frozen LLMs compared to fine-tuned counterparts. First, frozen LLMs may not effectively access the necessary knowledge to generate accurate responses. We demonstrate in Section 2.2 how to achieve this for the frozen LLM using probing techniques. Second, KGs often involve ambiguous entity representations, which can be difficult for LLM to handle accurately. In Section 2.3 we show how to achieve this with the generation of entity descriptions.
@@ 2.1 Formulation and Notations
A knowledge graph, $G$, comprises a set of entities, $E$, and relations, $R$. It can be represented as a set of triples $G=\{(h,r,t)\}$, where $h\in E$ is the head entity, $t\in E$ is the tail entity, we collectively refer to the head entity and tail entity as $e$, and $r\in R$ represents the relation between them. $D$ is the set of descriptions for each entity and relation. We denote $D(e)$, $D(r)$ as the textual description of each entity and each relation.
@@
2.2 Stimulating the Intermediate Layers of Large Language Models
Knowledge graphs consist of numerous triples, and the format of these triples is challenging for the LLM to understand, which may mislead the LLM to retrieve irrelevant knowledge and reveal the wrong responses. To enhance the performance of LLMs in KGC, we leverage the internal judgment mechanisms of the language model (Zou et al., 2023). Our method involves strategically triggering the necessary depth of knowledge embedded within the LLM to elicit the correct responses required for accurate KGC, thereby improving the overall performance of the KGC task.
Specifically, since different layers of the language model have different roles (Azaria and Mitchell, 2023), we determine the most suitable layer for the task by experimenting with the validation set. For the selection of intermediate layers, see Appendix B.5. First, we construct prompt templates to stimulate the model to generate the required classification information in the intermediate layers. In the triple classification, our goal is to determine whether the triples in the knowledge graph are true. Therefore, we use the following prompt template, $PT_{1}$, to create positive and negative sample pairs for training, as shown in Figure 2. Notice that the prompt does not describe positive or negative information for triples.
@@ 2.2 Stimulating the Intermediate Layers of Large Language Models
$v\left(s^{+}\right)=\{M_{1:i}\left(s^{+}\right)[-1],0<i<L\},$ (1)
where $M$ represents the language model, $M_{1:i}$ represents the first $i$ layers of the model, $[-1]$ refers to the hidden state corresponding to the last token, $L$ denotes the number of layers in the model. The dimension of $M_{1:i}(s^{+})[-1]$ is the embedding dimension of the language model. Negative samples follow the same principle.
At last, we train a classification model using $v(s^{+})$ and $v(s^{-})$ to distinguish between plausible and implausible triples. The goal is to learn a boundary that can separate the hidden states of the intermediate layers of the LLM corresponding to positive samples (which are true) from those corresponding to negative samples (which are false). We can use classification models such as multilayer perceptron (MLP), SVM, etc., for classification training. Taking logistic regression to determine the plausibility of samples as an example:
@@
where x represents $v(s)[l]$, $l$ is the layer index, $w$ and $b$ are weights and bias to be learned.
For relation prediction, we transform the binary classification into multi-class classification. By utilizing the prompt in Figure 7, we stimulate the LLM to generate hidden states that capture the characteristics of different relation types in the KG. Details can be found in the Appendix B.1.
By focusing on extracting classification results directly from the model’s internal processing, we ensure that the responses are both relevant and precise, thereby improving the effectiveness of the KGC task.
@@ 2.3 Subgraph Entity Description Generator
The semantics of entities in the knowledge graph are complex, making it challenging for the language model to interpret based solely on the entity’s name due to inherent ambiguity and polysemy. Therefore, we inform the language model about the meaning of each entity in the input. The meaning of each entity in KG is generated by its one-hop subgraph.
@@
To address incomplete entity descriptions within certain datasets, we have devised a method to generate $D(e)$. We construct $D(e)$ in two ways. The first approach involves a straightforward method where we concatenate all the triples within the one-hop subgraph surrounding the entity $e$, thereby forming $D(e)$ (Zhang et al., 2023). However, this direct concatenation can sometimes lead to ineffective or misleading stimuli for the LLM, as the structure of these subgraph triples may significantly differ from the text seen during the LLM’s pre-training phase. This divergence can adversely affect the model’s response accuracy. To mitigate this issue, our second approach leverages a language model to transform these triples into a format that is more comprehensible to the LLM. By rephrasing the triples from the one-hop subgraph around the entity $e$ into natural language text, we produce a more refined and model-friendly version of $D(e)$:
@@
where $\operatorname{subgraph(\cdot)}$ is a function that retrieves the set of all triples in the one-hop subgraph where the entity resides; $\operatorname{transform(\cdot)}$ is a function that converts the triple into sentence; $\operatorname{concat(\cdot)}$ is a function to process triple statements. For the first approach, $\operatorname{concat(\cdot)}$ is a string concatenation function that joins each transformed triple statement with a separator. For the second approach, we need to utilize LLM with the in-context learning technique (Brown et al., 2020) to generate descriptions of entities based on the triples in their subgraph. Please refer to Appendix A.1 for details.
@@ 3 Experiments
  • RQ3: What are the data efficiency and computation efficiency of our proposed method for KGC?
@@ 3.2 Experimental Settings
KG-LLAMA-SAMPLED N-L 3k/10k 0.727 0.682 0.614 0.574 0.834
LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.931 0.862
LLAMA-DES2 (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
LLAMA-DES2 (GPT)* H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
Table 2: The main experiment results (accuracy) of triple classification. In the $Pred$ column, $S\mbox{-}F$ uses score function, $L\mbox{-}MLP$ uses MLP on last layer hidden states, $N\mbox{-}L$ uses natural language, and $H\mbox{-}MLP$ uses MLP on intermediate layer hidden states to predict. $Samples$ indicates training sample counts, with $3k/10k$ indicating $10k$ samples for the UMLS dataset and $3k$ for others. LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimuli generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimuli generation. Top three results per dataset are bolded.
@@ 3.3 Main Results (For RQ1)
From Table 2, it can be observed that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples. Moreover, when trained on subsets of equivalent size, LLAMA-DES-H and LLAMA-DES2 (GPT) markedly surpass KG-LLAMA-SAMPLED, the existing state-of-the-art approach that relies exclusively on LLaMA for KGC. For more details, including the results of baseline models such as Alpaca-7B-ICL and Structural-aware-IT, and the performance of our method on the full training set, refer to Appendix B.4.
The experimental results presented in Table 2 also validate the effectiveness of the approaches proposed in Section 2.2 and Section 2.3. Specifically, LLAMA-DES-H and LLAMA-DES2 (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3. Furthermore, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B and LLAMA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2. Thus, the experimental results substantiate the effectiveness of our proposed methodology. Additionally, the frozen LLM can also achieve similar KGC effects as that of fine-tuning the LLM.
Notice that the performance on the UMLS dataset is special, where the use of generated descriptions hinders the model’s classification accuracy, and the use of non-generated entity descriptions is also worse than not using any entity descriptions. It is postulated that the underlying reason may lie in the highly specialized and domain-specific nature of entity names in the UMLS dataset, which are often precise medical terms with limited room for multiple interpretations. The entities in the knowledge graph have well-aligned with the internal entities of the LLM. Meanwhile, using entity descriptions may be affected by noisy description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states.
@@
3.4 Ablation Study (For RQ2)
Through the analysis depicted in the left portion of Figure 3, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline. Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers. This observation corroborates the notion posited by Geva et al. (2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers integrating information from lower layers. During pretraining, the lower layers may not have stored certain knowledge, and thus are unable to generate hidden states conducive to accurate triple classification. Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers. This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli (Jiang et al., 2024; Zou et al., 2023). Detailed explanations can be found in Appendix B.5. Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA. Meanwhile, as observed in Table 5, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
Regarding classification models, a comparative analysis encompassing MLP, SVM, and Logistic Regression, as delineated in Table 4, is conducted. The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP. This underscores the efficacy of our stimulus methodology, whereby hidden states of positive and negative samples within a high-dimensional space are easily distinguishable.